Drama made a pan of brownies for her family to share. On Monday, 3/4 of them were eaten. On Tuesday,1/2 of the remaining brownies were eaten. how much of the pan of brownies were left over for Wednesday?

Answer by algebra hello(55)

for tuesday the part of brownies left=1-3/4
=1/4
for wednesday the left over part of brownie=1/4-(1/4)(1/2)
=1/4-1/8
=1/8